# FlickerFusion: Intra-trajectory Domain Generalizing Multi-Agent RL

### v1

@@ Abstract
Multi-agent reinforcement learning has demonstrated significant potential in addressing complex cooperative tasks across various real-world applications. However, existing MARL approaches often rely on the restrictive assumption that the number of entities (e.g., agents, obstacles) remains constant between training and inference. This overlooks scenarios where entities are dynamically removed or added during the inference trajectory—a common occurrence in real-world environments like search and rescue missions and dynamic combat situations. In this paper, we tackle the challenge of intra-trajectory dynamic entity composition under zero-shot out-of-domain (OOD) generalization, where such dynamic changes cannot be anticipated beforehand. Our empirical studies reveal that existing MARL methods suffer significant performance degradation and increased uncertainty in these scenarios. In response, we propose FlickerFusion, a novel OOD generalization method that acts as a universally applicable augmentation technique for MARL backbone methods. Our results show that FlickerFusion not only achieves superior inference rewards but also uniquely reduces uncertainty vis-à-vis the backbone, compared to existing methods. For standardized evaluation, we introduce MPEv2, an enhanced version of Multi Particle Environments (MPE), consisting of 12 benchmarks. Benchmarks, implementations, and trained models are organized and open-sourced at flickerfusion305.github.io, accompanied by ample demo video renderings.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) is gaining significant attention in the research community as it addresses real-world cooperative problems such as autonomous driving (Antonio & Maria-Dolores, 2022), power grid control (Biagioni et al., 2022), and AutoML (Wang et al., 2023b). Nevertheless, MARL research is commonly over-simplified, in that often, restrictive assumptions are made. Among these assumptions, we focus on one where the number of entities (e.g., agents, obstacles) available during training and inference is the same (Wang et al., 2021b; Liu et al., 2023; Zang et al., 2023; Hu et al., 2024). In response, Iqbal et al. (2021) introduced the concept of dynamic team composition, where the number of agents varies between training and testing. This concept has been further explored in subsequent works (Hu et al., 2021; Wang et al., 2023a; Shao et al., 2023; Tian et al., 2023). Despite extensive research, two crucial characteristics ((1) and (2)) frequently encountered in real-world MARL deployments have been overlooked.
@@
These unexpected dynamic scenarios present a non-trivial problem—each agent is forced to observe an input-space greater than it was trained on. Prior works (Hu et al., 2021; Zhang et al., 2023; Shao et al., 2023) all took the approach of appropriately expanding the Q or policy network, via additional parameters, using inductive bias. However, empirically, all fail to deliver reasonable performance in the considered OOD MARL environment, let alone being consistent, across environments, and within environments (seeds)—even after extensive hyperparameter tuning (Table 1; Fig. 5 (top-left)). The high uncertainty of existing methods especially renders them unreliable for safety-critical applications (Knight, 2002) such as autonomous vehicles (Pek et al., 2020; Feng et al., 2023), where the deployed method must be robust to changes between train and test-time.
@@
We propose a novel, orthogonal approach for OOD MARL, FlickerFusion, that augments the agents’ observation space, introducing no additional parameters test-time. A high-level schematic overview is provided in Fig. 2. Prior works do not take this alternative approach as an immediate concern because, given no additional parameters, parts of the input space are lost. The stochasticity of the augmentation effectively preserves the information of the entire observation—aggregating the stochastic partial observations across time results in an emulated full observation.
@@
Note that the maximum number of entities need not be known beforehand, as the entity dropout can adapt to any OOD entity composition at test-time. The training is done similarly to “prepare” for the entity dropouts that the agents will experience at test time (Sec. 3). Remarkably, this simple lose and recover approach attains state-of-the-art performance vis-à-vis inference reward and uncertainty reduction, relative to an extensive list of existing MARL and other model-agnostic methods (Sec. 5). Finally, as no prior benchmarks can evaluate such dynamic MARL scenarios, we also standardize and present 12 benchmarks (Sec. 4).
@@ 2 Preliminaries
${\mathbb{N}}$ is the set of natural numbers, and ${\mathbb{N}}_{0}:=\{0\}\cup{\mathbb{N}}$. For a positive integer $n\in{\mathbb{N}}$, denote $[n]:=\{1,2,\cdots,n\}$. For two vectors ${\bm{a}}$ and ${\bm{b}}$ (possibly of different dimensions), ${\bm{a}}\oplus{\bm{b}}$ is the concatenated vector $[{\bm{a}}\ {\bm{b}}]$. For a set $A$, $|A|$ is its cardinality, $2^{A}$ is its power set, $\Delta(A)$ is the set of all possible probability distributions over $A$, whose support is potentially a strict subset of $A$, and $A^{*}$ is the set of all possible $A$-valued sequences of arbitrary lengths. Also, for an integer $n\leq|A|$, $\binom{A}{n}:=\{B\subset A:|B|=n\}$. For a $m\times n$ matrix ${\bm{A}}$, denote $\mathrm{nrow}({\bm{A}})=n$ to be the number of row.
@@
We follow the de facto MARL framework, Centralized Training Decentralized Execution (CTDE; Rashid et al. (2018); Foerster et al. (2017)). Our setting can be described as a Decentralized Markov Decision Process (Dec-MDP; Bernstein et al. (2002); Oliehoek & Amato (2016)) with dynamic entities (Iqbal et al., 2021; Schroeder de Witt et al., 2019), described by the tuple $\langle{\mathcal{E}}:={\mathcal{E}}_{a}\dot{\cup}{\mathcal{E}}_{n},\bm{\Phi},\phi,{\bm{S}},{\bm{U}},P,\widetilde{P},R,\gamma\rangle$.
Entities.
${\mathcal{E}}$ is the (potentially countably infinite) set of all possible entities, that may enter or leave intra-trajectory, and ${\mathcal{E}}$ is partitioned into the set of (trainable) agents ${\mathcal{E}}_{a}$ and other non-agent entities ${\mathcal{E}}_{n}$. For a $L\in{\mathbb{N}}$, we assume each entity takes one of $L$ types (e.g., ally, adversary, obstacle), and $\bm{\Phi}:=\{{\bm{e}}_{1},{\bm{e}}_{2},\cdots,{\bm{e}}_{L}\}$ is the set of all possible entity types where ${\bm{e}}_{\ell}$’s are the elementary bases of ${\mathbb{R}}^{L}$. For each entity $e\in{\mathcal{E}}$, we denote $\phi^{e}\in\Phi$ as its type.
@@
Transition Dynamics.
The dynamics consist of two parts: transitions in the entity composition and ${\bm{S}}$. Consider the current entity composition $E_{t}$ and state-action pair $({\bm{s}}_{t},{\bm{u}}_{t})\in{\bm{S}}(E_{t})\times{\bm{U}}(E_{t})$. The joint state transitions to ${\bm{s}}_{t+1}\sim P(\cdot|{\bm{s}}_{t},{\bm{u}}_{t},E_{t})\in\Delta({\bm{S}}(E_{t}))$; note how $P$ also depends on the current entity composition $E_{t}$. The entity composition transitions to $E_{t+1}\sim\widetilde{P}(\cdot|E_{t})\in\Delta(2^{{\mathcal{E}}})$.
@@ 3.1 Towards Domain Generalization Across Entity Composition
Figure 3: Example out-of-domain reward and uncertainty performance (color coded)
@@ 3.2 Flickering via Entity Dropout
Hereforth, while past works remedy OOD generalization via injecting inductive biases within $\bm{\theta}_{Q}^{\prime}$, our orthogonal approach ensures that $\mathrm{nrow}(\mathbf{X}^{\mathcal{D}_{O}})=\mathrm{nrow}(\mathbf{X}^{\mathcal{D}_{I}})$, no longer requiring $\bm{\theta}_{Q}^{\prime}$. In turn, the key design question becomes how many and which entities to drop.
@@
Now that we have determined how many entities of each type to drop, we discuss which ones to drop. Following Iqbal et al. (2021), we use randomized drop-outs: given $\Delta$, each agent independently and uniformly samples $\Delta[\ell]$ number of entities of type $\ell$ to drop. This simple solution is decentralized across the agents and, “on-average”, results in a “dispersed” view of the entities for the agents. This is formalized in the following statement, whose proof is deferred to Appendix A:
@@
$\mathbb{E}\left[\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\right]\leq\sqrt{\frac{\Delta_{\ell}(N_{\ell}-\Delta_{\ell})}{N_{A}}},$ (7)
which further implies the following high-probability guarantee: for any $\delta\in(0,1)$,
${\mathbb{P}}\left(\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\leq\frac{N_{\ell}}{N_{A}}\sqrt{\frac{1}{2}\log\frac{N_{\ell}}{\delta}}+\sqrt{\frac{\Delta_{\ell}(N_{\ell}-\Delta_{\ell})}{N_{A}}}\right)\geq 1-\delta.$ (8)
@@ 3.3 Fusing the Flickers
We now demonstrate how we can recover some of the lost information, further improving the trade-off of Eq. (6) to our favor. For each (active) agent $a$, let $D_{a}\subset E_{t}$ be the set of dropped entities from the perspective of $a$. If $D_{a}$ remains fixed throughout the trajectory, then $a$ never sees $D_{a}$. We alleviate this problem by stochastically changing $D_{a}$ at each $t$. In essence, FlickerFusion fuses (stochastic sampling at each $t$) the flickers (ED). The pseudocodes for train and inference modes are presented in Alg. 1 and 2. 1 Input: ${\bm{N}}^{train}$, $b$; 2 for each episode do 3       $\hat{{\bm{N}}}^{train}\leftarrow\text{Uniform}(\bm{1},{\bm{N}}^{train})$; 4       $\mathrm{CreateEnvironment}(\hat{{\bm{N}}}^{train})$; 5       for $t=1,2,\cdots$ do 6             $\Delta_{t}^{train}\leftarrow\text{Uniform}(\mathbf{0},\hat{{\bm{N}}}^{train})$; 7             ${\bm{o}}_{a},\_\leftarrow\text{ED}(a,\Delta_{t}^{train},E_{t})$ for each agent $a\in E_{t,a}$; 8             Add to central replay buffer; 9             if $t\ \%\ b==0$ then 10                   Batch sample from replay buffer; 11                   Train ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\theta}}$ via QMIX; 12             13       Algorithm 1 FlickerFusion (Train) 1 Input: ${\bm{N}}^{train}$; 2 for $t=1,2,\cdots$ do 3       for $a\in E_{t,a}$ (decentrally) do 4             if $a$ is newly introduced then 5                   Initialize $\tau^{a}\leftarrow\{\}$; 6                   7            $\Delta_{t}^{inf}\leftarrow\text{Observe}(t,a)-{\bm{N}}^{train}$; 8             ${\bm{o}}_{a},N_{t,a}\leftarrow\text{ED}(a,\Delta_{t}^{inf},E_{t})$; 9             $\tau^{a}\leftarrow\tau^{a}\cup\{{\bm{o}}_{a}\}$; 10             Choose $u_{a}\leftarrow\operatorname*{arg\,max}_{u}Q^{a}(\tau^{a},u;\bm{\theta}_{Q}^{a}\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{\theta}})$; 11             $\tau^{a}\leftarrow\tau^{a}\cup\{\{u_{e}\}_{e\in N_{t,a}}\}$; 12             13       Algorithm 2 FlickerFusion (Inference)
@@
Alg. 2 is the pseudocode for FlickerFusion in the inference phase, again taking the in-domain entity composition ${\bm{N}}^{train}$ as input. At each time $t$, the number of entities to be dropped $\Delta_{t}^{inf}$ is computed from the perspective of agent $a$, where $\text{Observe}(t,a)$ outputs the entity combination vector observed by agent $a$ (line 6). We implement stochastically changing $D_{a}$ by
@@
calling ED at each $t$ such that the random partial observation (${\bm{o}}_{a}$), as well as the set of agents not dropped ($N_{t,a}$), are obtained from DAED (line 7), using which the next action is chosen greedily (line 9). Importantly, all this is done decentrally across agents at each time $t$.
To concretely demonstrate the effect of calling DAED every $t$ (line 7), consider a toy scenario where $L=1$, $\mathrm{nrow}({\bm{X}}^{{\mathcal{D}}_{I}})=4$, and $\mathrm{nrow}({\bm{X}}^{{\mathcal{D}}_{O}})=5$. As illustrated in Fig. 4, each agent’s view stochastically changes at each $t$. The important intuition is that aggregating these views along the temporal axis gives us a virtually full view; this is illustrated more clearly on our project site.

### v4

@@ Abstract
Multi-agent reinforcement learning has demonstrated significant potential in addressing complex cooperative tasks across various real-world applications. However, existing MARL approaches often rely on the restrictive assumption that the number of entities (e.g., agents, obstacles) remains constant between training and inference. This overlooks scenarios where entities are dynamically removed or added during the inference trajectory—a common occurrence in real-world environments like search and rescue missions and dynamic combat situations. In this paper, we tackle the challenge of intra-trajectory dynamic entity composition under zero-shot out-of-domain (OOD) generalization, where such dynamic changes cannot be anticipated beforehand. Our empirical studies reveal that existing MARL methods suffer significant performance degradation and increased uncertainty in these scenarios. In response, we propose FlickerFusion, a novel OOD generalization method that acts as a universally applicable augmentation technique for MARL backbone methods. FlickerFusion stochastically drops out parts of the observation space, emulating being in-domain when inferenced OOD. The results show that FlickerFusion not only achieves superior inference rewards but also uniquely reduces uncertainty vis-à-vis the backbone, compared to existing methods. Benchmarks, implementations, and model weights are organized and open-sourced at flickerfusion305.github.io, accompanied by ample demo video renderings.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) is gaining significant attention in the research community as it addresses real-world cooperative problems such as autonomous driving (Antonio & Maria-Dolores, 2022), power grid control (Biagioni et al., 2022), and AutoML (Wang et al., 2023b). Nevertheless, MARL research is commonly over-simplified, in that often, restrictive assumptions are made. Among these assumptions, we focus on one where the number of entities (e.g., agents, obstacles) available during training and inference is the same (Wang et al., 2021b; Liu et al., 2023; Zang et al., 2023; Hu et al., 2024). In response, Iqbal et al. (2021) introduced the concept of dynamic team composition, where the number of agents varies between training and testing. This concept has been further explored in subsequent works (Hu et al., 2021; Wang et al., 2023a; Shao et al., 2023; Tian et al., 2023; Yuan et al., 2024). Other variations including situations where the policies of uncontrollable teammates arbitrarily change intra-trajectory have also been studied (Zhang et al., 2023b). Despite extensive research, two crucial characteristics ((1) and (2)) frequently encountered in real-world MARL deployments have been overlooked.
@@
These unexpected dynamic scenarios present a non-trivial problem—each agent is forced to observe an input-space greater than it was trained on. Prior works (Hu et al., 2021; Zhang et al., 2023a; Shao et al., 2023) all took the approach of appropriately expanding the Q or policy network, via additional parameters, using inductive bias. However, empirically, all fail to deliver reasonable performance in the considered OOD MARL environments, let alone being consistent, across environments, and within environments (seeds)—even after extensive hyperparameter tuning (Table 1; Fig. 5 (top-left)). The high uncertainty of existing methods especially renders them unreliable for safety-critical applications (Knight, 2002) such as autonomous vehicles (Pek et al., 2020; Feng et al., 2023), where the deployed method must be robust to changes between train and test-time.
@@
We propose a novel, orthogonal approach for OOD MARL, FlickerFusion, that augments the agents’ observation space, introducing no additional parameters test-time. The augmentation is based on the widely-adopted dropout mechanism (Srivastava et al., 2014)—where neurons (including the input layer), are randomly masked out. FlickerFusion differs from Srivastava et al. (2014) as it leverages the dropout mechanism only in the input layer, such that, the model can remain in an emulated in-domain regime, even when inferenced OOD. A high-level schematic overview is provided in Fig. 2. Prior works do not take this alternative approach as an immediate concern becomes, given no additional parameters, parts of the input space are lost.
@@
The stochasticity of the augmentation effectively preserves the information of the entire observation—aggregating the stochastic partial observations across time results in an emulated full observation. Note that the maximum number of entities need not be known beforehand, as the entity dropout can adapt to any OOD entity composition at test-time. The training is done similarly to “prepare” for the entity dropouts that the agents will experience at test time (Sec. 3). Remarkably, this simple lose and recover approach attains state-of-the-art performance vis-à-vis inference reward and uncertainty reduction, relative to an extensive list of existing MARL and other model-agnostic methods (Sec. 5). Finally, as no prior benchmarks can evaluate such dynamic MARL scenarios, we also standardize and present 12 benchmarks (Sec. 4).
@@ 2 Preliminaries
For a positive integer $n\in{\mathbb{N}}$, denote $[n]:=\{1,2,\cdots,n\}$. For two vectors ${\bm{a}}$ and ${\bm{b}}$, ${\bm{a}}\oplus{\bm{b}}$ is the concatenated vector $[{\bm{a}}\ {\bm{b}}]$. For a set $A$, $|A|$ is its cardinality, $2^{A}$ is its power set, $\Delta(A)$ is the probability distribution’s possibility set over $A$, and $A^{*}$ is the set of all possible $A$-valued sequences of arbitrary lengths. Also, for an integer $n\leq|A|$, $\binom{A}{n}:=\{B\subset A:|B|=n\}$. For a $m\times n$ matrix ${\bm{A}}$, denote $\mathrm{nrow}({\bm{A}})=n$ to be the number of row.
@@
We follow the de facto MARL framework, Centralized Training Decentralized Execution (CTDE; Rashid et al. (2018); Foerster et al. (2017)). Our setting can be described as a Decentralized Markov Decision Process (Dec-MDP; Bernstein et al. (2002); Oliehoek & Amato (2016)) with dynamic entities (Iqbal et al., 2021; Schroeder de Witt et al., 2019), described by the tuple $\langle{\mathcal{E}}:={\mathcal{E}}_{a}\dot{\cup}{\mathcal{E}}_{n},\bm{\Phi},\phi,{\bm{S}},{\bm{U}},P,\widetilde{P},R,\gamma\rangle$. ${\mathcal{E}}$ is the set of all possible entities, that may enter or leave intra-trajectory, and ${\mathcal{E}}$ is partitioned into the set of (trainable) agents ${\mathcal{E}}_{a}$ and other non-agent entities ${\mathcal{E}}_{n}$. For a $L\in{\mathbb{N}}$, we assume each entity takes one of $L$ types (e.g., ally, adversary, obstacle), and $\bm{\Phi}:=\{{\bm{e}}_{1},{\bm{e}}_{2},\cdots,{\bm{e}}_{L}\}$ is the set of all possible entity types where ${\bm{e}}_{\ell}$’s are the elementary bases of ${\mathbb{R}}^{L}$. For each entity $e\in{\mathcal{E}}$, we denote $\phi^{e}\in\Phi$ as its type.
@@ 3.1 Towards Domain Generalization Across Entity Composition
Figure 3: Example out-of-domain reward and uncertainty performance (color coded). Shaded region in (b) indicate quartile 1 to 3.
@@ 3.2 Flickering via Entity Dropout
While past works remedy OOD generalization via injecting inductive biases within $\bm{\theta}_{Q}^{\prime}$, our orthogonal approach ensures that $\mathrm{nrow}(\mathbf{X}^{\mathcal{D}_{O}})=\mathrm{nrow}(\mathbf{X}^{\mathcal{D}_{I}})$, no longer requiring $\bm{\theta}_{Q}^{\prime}$. In turn, the key design question becomes how many and which entities to drop.
@@
Now that we have determined how many entities of each type to drop, we discuss which ones to drop. Following Iqbal et al. (2021), we use randomized drop-outs: given $\Delta:={\bm{N}}^{inf}-{\bm{N}}^{train}$, each agent independently and uniformly samples $\Delta[\ell]$ number of entities of type $\ell$ to drop. This simple solution is decentralized across the agents and, “on-average”, results in a “dispersed” view of the entities for the agents. This is formalized in the following statement, whose proof is deferred to Appendix A:
@@
$\mathbb{E}\left[\sum_{i\in[N_{\ell}]}\left|\frac{d(i)}{N_{A}}-\frac{\Delta_{\ell}}{N_{\ell}}\right|\right]\leq\sqrt{\frac{\Delta_{\ell}(N_{\ell}-\Delta_{\ell})}{N_{A}}}.$ (7)
@@ 3.3 Fusing the Flickers
We now demonstrate how we can recover some of the lost information, further improving the trade-off of Eq. (6) to our favor. For each (active) agent $a$, let $D_{a}\subset E_{t}$ be the set of dropped entities from the perspective of $a$. If $D_{a}$ remains fixed throughout the trajectory, then $a$ never sees $D_{a}$. We alleviate this problem by stochastically changing $D_{a}$ at each $t$. This also has the added benefit that any intra-trajectory dynamics changes can be immediately responded to on-the-fly. In essence, FlickerFusion fuses (stochastic sampling at each $t$) the flickers (ED). The pseudocodes for train and inference modes are presented in Alg. 1 and 2.
@@
Alg. 2 is the pseudocode for FlickerFusion in the inference phase, again taking the in-domain entity composition ${\bm{N}}^{train}$ as input. At each time $t$, the number of entities to be dropped $\Delta_{t}^{inf}$ is computed from the perspective of agent $a$, where $\text{Observe}(t,a)$ outputs the entity combination vector observed by agent $a$ (line 6). We implement stochastically changing $D_{a}$ by calling ED at each $t$ such that the random partial observation (${\bm{o}}_{a}$), as well as the set of agents not dropped ($N_{t,a}$), are obtained from DAED (line 7), using which the next action is chosen greedily (line 9). Importantly, all this is done decentrally across agents at each time $t$. 1 Input: ${\bm{N}}^{train}$, $b$; 2 for each episode do 3       $\hat{{\bm{N}}}^{train}\leftarrow\text{Uniform}(\bm{1},{\bm{N}}^{train})$; 4       $\mathrm{CreateEnvironment}(\hat{{\bm{N}}}^{train})$; 5       for $t=1,2,\cdots$ do 6             $\Delta_{t}^{train}\leftarrow\text{Uniform}(\mathbf{0},\hat{{\bm{N}}}^{train})$; 7             ${\bm{o}}_{a},\_\leftarrow\text{ED}(a,\Delta_{t}^{train},E_{t})$ for each agent $a\in E_{t,a}$; 8             Add to central replay buffer; 9             if $t\ \%\ b==0$ then 10                   Batch sample from replay buffer; 11                   Train ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{\theta}}$ via QMIX; 12             13       Algorithm 1 FlickerFusion (Train) 1 Input: ${\bm{N}}^{train}$; 2 for $t=1,2,\cdots$ do 3       for $a\in E_{t,a}$ (decentrally) do 4             if $a$ is newly introduced then 5                   Initialize $\tau^{a}\leftarrow\{\}$; 6                   7            $\Delta_{t}^{inf}\leftarrow\text{Observe}(t,a)-{\bm{N}}^{train}$; 8             ${\bm{o}}_{a},N_{t,a}\leftarrow\text{ED}(a,\Delta_{t}^{inf},E_{t})$; 9             $\tau^{a}\leftarrow\tau^{a}\cup\{{\bm{o}}_{a}\}$; 10             Choose $u_{a}\leftarrow\operatorname*{arg\,max}_{u}Q^{a}(\tau^{a},u;\bm{\theta}_{Q}^{a}\in{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{\theta}})$; 11             $\tau^{a}\leftarrow\tau^{a}\cup\{\{u_{e}\}_{e\in N_{t,a}}\}$; 12             13       Algorithm 2 FlickerFusion (Inference)
@@
To concretely demonstrate the effect of calling DAED every $t$ (line 7), consider a toy scenario where $L=1$, $\mathrm{nrow}({\bm{X}}^{{\mathcal{D}}_{I}})=4$, and $\mathrm{nrow}({\bm{X}}^{{\mathcal{D}}_{O}})=5$. As illustrated in Fig. 4, each agent’s view stochastically changes at each $t$. As seen by the green, executing the algorithm at every $t$ ensures smooth transition to any intra-trajectory OOD dynamics. The important intuition is that aggregating these views along the temporal axis gives us a virtually full view; this is illustrated more clearly on our project site.